Suppose that in a recent market period, the following relationship existed between the price of blood glucose meters and the quantity supplied and quantity demanded:

Quantity Demanded

Price

Quantity Supplied

0

$50

80,000

10,000

$45

70,000

20,000

$40

60,000

30,000

$35

50,000

40,000

$30

40,000

50,000

$25

30,000

60,000

$20

20,000

70,000

$15

10,000

80,000

$10

0

A. What are the equilibrium price and quantity?

B. If the industry price is $25, is there a shortage or surplus of blood glucose meters? Please explain.

C. If the industry price is $35, is there a shortage or surplus of blood glucose meters? Please explain.   
Answer #1

Answer
A)
the equilibrium is at
quantity demanded =quantity supplied
it is at Q=40000
where
P=$30
the equilibrium price is $30 and quantity is 40000 units
=======
B)
P=$25
Qd=50000 and Qs=30000
Qd>Qs means there is a shortage, as the quantity demanded is higher than the quantity supplied.
shortage =Qd-Qs=50000-30000=20000
the shortage is 20000 units

C)
P=35
Qd=30000 and Qs=50000
Qs>Qd means there is a surplus of goods in the market as the quantity supplied is higher than the quantity demanded
surplus =Qs-Qd=50000-30000=20000
the surplus is 20000 units